OpenStudy (anonymous):

Okay, how do I convert from standard form of a circle to general form? Specifically, the equation I'm currently trying to convert is (x + 4)^2 + (y - 2)^2 = 4

OpenStudy (anonymous):

what is a general form?

OpenStudy (anonymous):

x^2 + y^2 + ax + by + c = 0

OpenStudy (anonymous):

^ general form

zepdrix (zepdrix):

Hmm so I guess you just have to expand out the brackets. Do you remember how to multiply a pair of binomials?\[\Large\bf\sf (x+4)^2\quad=\quad (x+4)(x+4)\quad=\quad ?\]

OpenStudy (anonymous):

Hmm, yes, hold on let me try this, I think I might know it. (x + 4)^2 = x^2 + 8x + 16 (y - 2)^2 = y^2 - 4y + 4 So x^2 + y^2 + 8x - 4y + 20 = 4 x^2 + y^2 + 8x - 4y + 16 = 0

OpenStudy (anonymous):

Is that right? Did I do it right?

zepdrix (zepdrix):

Hmm why did your +4 from the y brackets get moved to the right side? :o

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

It wasn't, but in the standard form, the equation was equal to 4, so I had to move that over to the left side to get 0

zepdrix (zepdrix):

Ohhhh it started equal to 4 haha :) good good good. Ok yes good job!

OpenStudy (anonymous):

Yaaay! lol First time I've ever figured something like that out xD thank you!

OpenStudy (anonymous):

you're welcome :)